OpenStudy (anonymous):

WILL GIVE MEDAL! Match each actual value with the predicted value that results in a thirty−five percent error. Actual: two hundred Actual: one hundred eighty Actual: one hundred sixty Actual: two hundred twenty A) Predicted: two hundred forty−three B) Predicted: two hundred sixteen C) Predicted: two hundred seventy D) Predicted: one hundred forty−three

OpenStudy (anonymous):

@greenlegodude57

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

(actual - predicted)/(actual) = 0.3

OpenStudy (anonymous):

?

ganeshie8 (ganeshie8):

or, (actual - predicted)/(actual) = -0.3

ganeshie8 (ganeshie8):

oh wait a sec, the error percent is 35 so, the predicted can be :- solving both gives u :- predicted = 0.65(actual) or predicted = 1.35(actual)

ganeshie8 (ganeshie8):

lets take := Actual: two hundred

ganeshie8 (ganeshie8):

predicted can be :- 0.65(200) = 130 or 1.35(200) = 270

ganeshie8 (ganeshie8):

since u have 270 on the other side, pick that one for Actual: two hundred

ganeshie8 (ganeshie8):

Actual: two hundred (C) Actual: one hundred eighty Actual: one hundred sixty Actual: two hundred twenty

ganeshie8 (ganeshie8):

see if u can figure out remaining same way ?

OpenStudy (anonymous):

But how do you get decimals?

ganeshie8 (ganeshie8):

wat decimals ?

OpenStudy (anonymous):

You did this above: 1.35(200) = 270

OpenStudy (anonymous):

How did you get the 1.35?

ganeshie8 (ganeshie8):

good question :) we're given that the error is 35% right ?

ganeshie8 (ganeshie8):

so, the predicted value will be either : (100+35)% or (100-35)% of the actual value

ganeshie8 (ganeshie8):

that gives u 135% or 65 %

ganeshie8 (ganeshie8):

which is same as 1.35 or 0.65

OpenStudy (anonymous):

Ok.

ganeshie8 (ganeshie8):

see if u can match the table

OpenStudy (anonymous):

Wait, so we add one of the actual value with 0.35?

ganeshie8 (ganeshie8):

yup! or subtract 35%

OpenStudy (anonymous):

The actual 200 is the predicted 270.

OpenStudy (anonymous):

The actual 180 is the predicted 243.

ganeshie8 (ganeshie8):

yup ! 270 = 200 + 35% of 200

OpenStudy (anonymous):

The actual 160 is the predicted 216.

ganeshie8 (ganeshie8):

You got it !!!!

OpenStudy (anonymous):

And the actual 220 is the predicted 143. We had to subtract there.

OpenStudy (anonymous):

Wow. Thanks so much @ganeshie8 !

ganeshie8 (ganeshie8):

exactly ! good job :)

OpenStudy (anonymous):

Have a good medal! :)

ganeshie8 (ganeshie8):

u deserve it :D

OpenStudy (anonymous):

No, you do!